OpenStudy (anonymous):

This is the last one for now. I hope you guys are on when I take my quiz because there are two problems like this on that I can't get. 1/9x^2+11/3x=432

OpenStudy (anonymous):

This is just like the last problem except where there was 80 you put 432

OpenStudy (anonymous):

multiply every term by x^2 and then solve using the quadratic formula.

myininaya (myininaya):

i think the x^2 is suppose to be in the numerator

OpenStudy (anonymous):

adfriedm I believe the x^2 is on top... just guessing.

OpenStudy (anonymous):

fair enough

OpenStudy (anonymous):

I don't get the quadratic formula. To me it doesn't make sense

OpenStudy (anonymous):

Blasphemy

myininaya (myininaya):

you can find that formula if you complete the square for ax^2+bx+c=0 so it makes perfect sense

OpenStudy (anonymous):

is it (1/9)x^2+(11/3)x=432 or 1/(9x^2)+11/(3x)=432?

OpenStudy (anonymous):

of course it requires a to be non-zero

myininaya (myininaya):

rihgt! because the quadratic formula is undefined for a=0

OpenStudy (anonymous):

well moreover because the proof doesn't follow with a=0

OpenStudy (anonymous):

You're just arguing semantics...

OpenStudy (anonymous):

not in the slightest

OpenStudy (anonymous):

I think what confuses me and I know this is stupid. What is the + sign with the line under it?

myininaya (myininaya):

it means you have 2 annswers one with + and one with minus

OpenStudy (anonymous):

Ok but how do I work this if I have a + and - answer. I'm sorry I just don't get it

myininaya (myininaya):

\[{-b+\sqrt{b^2-4ac}}/(2a) and ( -b-\sqrt{b^2-4ac})/(2a)\]

OpenStudy (anonymous):

it is the a quadratic equation, so we always expect 2 values which make it zero.

OpenStudy (anonymous):

OH ok. You guys are a life saver!!! thank you

OpenStudy (anonymous):

\[\frac{1}{9 x^2}+\frac{11}{3 x}=432 \] Put the 432 on the left hand side of the equation and combine terms into one fraction. \[\frac{1+33 x-3888 x^2}{9 x^2}=0 \] If the fraction is zero, the Numerator is zero. Factor the Numerator and set the result to zero. x = 1/48 and x = -1/81

OpenStudy (anonymous):

Thank you